# FreeInsert: Disentangled Text-Guided Object Insertion in 3D Gaussian Scene without Spatial Priors

### v2

@@ 3.3. Initialization via Large Models
To initialize the $\textit{t}_{\textit{O}}$, we use strong pixel-level semantic spatial localization capability of Molmo (Deitke et al., 2024) to predict a set of 2D object centers $\{c_{O}^{(v)}\}_{v\in\mathcal{V}}$ across multiple scene views with the Local Target $\mathcal{T}_{\textit{LT}}$, using prompt like “Point the position to add ¡$\mathcal{T}_{\textit{LT}}$¿”. Let $\hat{\mathcal{G}}_{O}^{(t)}$ denote the object geometry after applying the transformation $(\textit{s}_{\textit{O}}$, $\textit{r}_{\textit{O}}$, $t$ ), where $t$ is a optimized parameter. For each view $v$, we project the transformed object and compute the 2D centroid of its projection. The $\textit{t}_{\textit{O}}$ is obtained by optimizing $t$ to minimize the discrepancy between the projected and the predicted centroids:
@@ 4. Experiments
Implementation Details During initialization, we use a learning rate of $5\times 10^{-3}$ for optimizing both $\mathcal{G}_{\textit{AR}}$ and $\textit{t}_{\textit{O}}$ of inserted object. When estimating the coarse rotation $\textit{r}_{\textit{O}}$, we render the object at 10-degree intervals . During the Hierarchical Spatial Aware Refinement stage, we apply a learning rate of $5\times 10^{-4}$ with diffusion timesteps in the range of [0.02, 0.2], $\lambda=0.1$ is set in $\mathcal{L}_{loc}$ and $\beta$ is linearly increased from 0 to 1 during training. For appearance refinement, we optimize the object appearance using timesteps in [0.02, 0.5$\sim$0.25]. The object image $\mathcal{I}_{\textit{O}}$ is upsampled with a sampling ratio of $M/N=3$ relative to multi-view inputs. All experiments are conducted on a single NVIDIA A40 GPU. More details are provided in the Appendix.
@@ 4.3. Ablation Studies
Effectiveness of Viewpoint Frequency Balancing. Figure 10 compares object appearance optimization when fine-tuning LoRA with a single object image versus combining it with multi-view images at different sampling frequency ratios (${FR}$). Using only inserted-object image leads to shape inconsistencies and artifacts across views (item (b)), while incorporating multi-view data improves consistency but may reduce detail. Our experiments show that ${FR=3}$ offers the best trade-off between multi-view consistency and single-view image quality.

### v3

@@ 3.3. Initialization via Large Models
To initialize the $\textit{t}_{\textit{O}}$, we use strong pixel-level semantic spatial localization capability of Molmo (Deitke et al., 2024) to predict a set of 2D object centers $\{c_{O}^{(v)}\}_{v\in\mathcal{V}}$ across multiple scene views with the Local Target $\mathcal{T}_{\textit{LT}}$, using prompt like “Point the position to add $\mathcal{T}_{\textit{LT}}$¿”. Let $\hat{\mathcal{G}}_{O}^{(t)}$ denote the object geometry after applying the transformation $(\textit{s}_{\textit{O}}$, $\textit{r}_{\textit{O}}$, $t$ ), where $t$ is a optimized parameter. For each view $v$, we project the transformed object and compute the 2D centroid of its projection. The $\textit{t}_{\textit{O}}$ is obtained by optimizing $t$ to minimize the discrepancy between the projected and the predicted centroids:
@@ 4. Experiments
[…]
@@ 4.3. Ablation Studies
Effectiveness of Viewpoint Frequency Balancing. Figure 10 compares object appearance optimization when fine-tuning LoRA with a single object image versus combining it with multi-view images at different sampling frequency ratios (FR{FR}). Using only inserted-object image leads to shape inconsistencies and artifacts across views (item (b)), while incorporating multi-view data improves consistency but may reduce detail. Our experiments show that FR=3{FR=3} offers the best trade-off between multi-view consistency and single-view image quality.